# DNN Task Assignment in UAV Networks: A Generative AI Enhanced Multi-Agent Reinforcement Learning Approach

### v2

@@ Abstract
Unmanned Aerial Vehicles (UAVs) offer high mobility and flexible deployment capabilities, making them ideal for Internet of Things (IoT) applications. However, the substantial amount of data generated by various applications within the existing low-altitude economic network requires processing through Deep Neural Networks (DNN) on UAVs, which is challenging due to their limited computational resources. To address this issue, we propose a two-stage optimization method for flight path planning and task allocation based on a mother-child UAV swarm system. In the first stage, we employ a greedy algorithm to solve the path planning problem by considering the task size of the target area to be inspected and the shortest flying path as constraints. The goal is to minimize both the UAV’s flight path and the overall system cost. In the second stage, we introduce a novel DNN task assignment algorithm that combines Multi-Agent Deep Deterministic Policy Gradient (MADDPG) and Generative Diffusion Models (GDM), named GDM-MADDPG. This algorithm leverages the reverse denoising process of GDM to replace the actor network in MADDPG. It enables UAVs to generate specific DNN task assignment actions based on agents’ observations in a dynamic environment, thereby improving the efficiency of task scheduling and overall system performance. Simulation results demonstrate that our algorithm outperforms benchmarks in terms of path planning, Age of Information (AoI), task completion, and system utility, showcasing its effectiveness and practical advantages in real-world scenarios.
@@ I-A Background Research
Advancements in Unmanned Aerial Vehicles (UAVs) technology have led to their increasing use in various civil and military domains, including aerial object detection, rapid rescue operations in disaster-stricken areas [1, 2], emergency scenarios [3], and extensive agriculture [4]. Unlike traditional Internet of Things (IoT) devices [5], UAVs offer not only lower costs, but also superior mobility and enhanced data collection capabilities [6]. Although UAVs are well suited for the critical and complex tasks mentioned, managing the substantial data they generate remains a challenge [7]. With the rise of artificial intelligence (AI), deep neural networks (DNN) have been proposed to handle the extensive data produced [8]. Despite their efficiency in processing data, DNNs require significant resources due to their intricate structure and high performance demands. General methodologies segmented DNN models by layers, treating each layer as the primary computational unit, which often led to inefficiencies in collaborative inference due to the substantial computational demands of certain layers. Given the size and cost constraints of UAVs, their power supply is often limited and computing capacity is relatively low. Generally, there are two primary approaches to tackle this issue. One approach involves deploying lightweight models on UAVs, which reduces computational demands and inference latency but sacrifices accuracy [9]. However, executing computation-intensive tasks independently with a single UAV presents significant challenges in achieving low latency and high inference accuracy. Another approach is to design from the perspective of the end-edge-cloud network architecture, enabling real-time, high-accuracy processing on UAVs with support from cloud and edge servers.
@@
Ref. Architecture Adaptive partitioning Path planning Latency Energy consumption Reliability Age of information
@@ I-B Motivation and Challenges
Figure 2: The computational complexity, processing latency, and output data size of each layer of different types of DNNs.
@@ I-C Summary of Contributions
  • We propose a mother-and-child UAV system that employs a high-altitude platform (HAP), such as an airship, equipped with multiple UAV launchers. From this HAP, several UAVs are launched and networked to support collaborative tasks. Then we design DNN tasks directly on UAVs, performing inference independently of the ground base station. This approach enhances the scalability of the system. Furthermore, we propose a novel greedy algorithm for UAV path planning that utilizes a fitness function to minimize flying distance, taking into account the varying sizes of tasks.
@@
Related Works
Henceforth, we summarize contributions of the related works and discuss the distinctions between prior research and our methodology.
@@ II-A Edge Intelligent for DNN Task
The partition of the DNN task represents a significant research challenge that involves dividing a DNN into several segments and distributing these segments between designated locations. The existing work provides valuable information to support the efficient DNN inference task. The authors in [24] present a joint design for task partitioning and offloading in DNN-enabled edge computing, involving a server and multiple devices. In [25], the authors propose a multi-task learning-based asynchronous advantage actor-critic approach, and explore distributed DNN inference through fine-grained model partitioning in edge computing. These studies overlook the mobility of UAV networks. The authors in [26] explore accelerating DNN-based task processing in vehicular edge computing through task partitioning and offloading. In [17], the authors propose an efficient allocation strategy using RL and investigate fast DNN inference in UAV swarms through multi-UAV collaboration. Studies have not addressed the adaptation challenges associated with the partitioning of DNN tasks. However, the oversight of these studies may result in decreased network reliability in certain resource-constrained networks due to suboptimal task assignment within the DNN.
@@ II-B MARL for Task Offloading
MARL enables each agent to optimize its policy through observations of both the environment and the policies of other agents [27, 28]. Many existing MARL-based approaches significantly improve task offload efficiency within dynamic networks [29, 30]. The authors in [31] propose a dedicated actor neural network for coordination and a scalable training algorithm for UAV-aided MEC task offloading. In [32], the authors present a DRL-based joint safe offloading and resource allocation scheme for vehicular edge computing networks. These studies have paid limited attention to this aspect. The authors in [33] present a mobility-aware service offloading and migration scheme with Lyapunov optimization and a MADDPG algorithm. The authors in [34] focus on optimizing task offloading and UAV trajectory under energy and queue latency constraints. However, most of these studies focus on optimizing network performance metrics, including task completion latency and energy consumption. Our research emphasizes DNN inference services, where the AoI for DNN tasks is a critical performance metric.
@@ II-C GAI for UAV Networks
GAI has the potential to effectively overcome the limitations of conventional AI and can be applied to optimize UAV networks, particularly to improve transmission rates, communication capacities, and energy efficiency [35, 36]. In [37], the authors present a radio propagation model with a conditional generative adversarial network. The authors in [38] introduce a two-stage generative neural network to predict link states and generate path losses, latencys, and arrival and departure angles in millimeter wave communication for UAVs. The authors in [39] explore the long-term optimization of joint task offloading and resource allocation in a multi-access edge computing network for UAVs. In [40], the authors present a hierarchical aerial computing system where UAVs collect tasks from ground devices and offload them to HAPs. Unlike the above schemes, we contribute to the existing literature by presenting a novel application of GDM in the context of DNN task assignment decision. Our approach incorporates a path planning, GDM-based MARL decision-making framework specifically tailored for multi-UAV networks. This methodology aims to minimize the cost associated with DNN task completion while ensuring the system’s reliability.
System Model
In this section, we first present the considered network architecture, then introduce in detail the DNN task model, the communication model, the mobile model, the AoI model, and the energy consumption model, respectively.
@@ III-A Network Architecture
The network architecture is primarily composed of an HAP layer and a UAV swarm layer, as illustrated in Fig. 3. The HAP layer consists of a tethered platform with extensive computing and communication capabilities, such as an airship. This platform exchanges information with UAVs via wireless links to facilitate model training [41] and path planning (see Section V). The airship will transmit the final result of the path planning to the leader UAV. The UAV swarm comprises a leader UAV and multiple wingman UAVs. All UAVs are isomorphic, with the total number of UAVs being ${n}=\left\{{1,2,\cdot\cdot\cdot,N}\right\}$, and each is equipped with trained DNN models. The leader UAV receives the results of the path planning transmitted by the airship and operates according to the predetermined route. It is also responsible for collecting and processing images and point cloud data of the target area, including tasks such as target detection and mapping. Thus, it functions as both a task producer and executor, as well as a decision maker for task assignment. The wingman UAVs primarily execute the tasks offloaded by the leader UAV, with the assignment of DNN tasks between wingman UAVs being streamlined. The leader UAV evaluates the task size and the remaining resource status of the wingman UAVs to determine whether to offload tasks. Therefore, to prevent task overload on the UAVs, extend the operational lifespan of the UAV swarm, and improve the timeliness of task processing and network stability, efficient task assignment decisions for the UAV swarm are particularly crucial.
@@ III-B DNN Task Model
The area where the task is performed in this document consists of $q$ center coordinate points of the target area, denoted as $q=1,2,...,\mathbb{W}$, and the sequence of the UAV inspection target points is represented as $\mathfrak{q}(q)=[\mathfrak{q}(1),\ldots,\mathfrak{q}(\mathbb{W})]$. The data size in the target area is $\ell_{q}=\{1,...,\mathbb{L}_{\mathbb{W}}\}$. Assume that the number of DNN tasks in each synchronization cycle is $i=\{1,2,...,I\}$, the DNN type is $\kappa=\{1,...,\mathbb{K}\}$. The DNN model typically comprises $l$ layers, denoted as $l=\{1,2,...,L\}$. Each layer has a cache capacity $m_{i,l}$ and a computing requirement ${c}_{i,l}$, represented by a tuple $(m_{i,l},c_{i,l})$. $l_{\kappa}=\{1,2,...,L_{\kappa}\}$ represents a set of layers associated with a specific type of DNN task, including convolutional and pooling layers for feature extraction, as well as fully connected layers for classification [8]. The DNN task is decomposed into multiple subtasks for serial processing, and each wingman UAV serves as an execution stage in this serial processing. This approach allows each subtask to offload and transfer intermediate results among different wingman UAVs without requiring the processing of the entire task, thereby enhancing the computational efficiency and real-time performance of computationally intensive and latency-sensitive tasks [11].
In response to such practical needs, determining the optimal split point of the task is crucial in DNN task assignment. Assume that the split point of the ${i}$-th DNN task model is $P_{i,n^{\prime}}$, and $n^{\prime}\in\{1,2,...,N-1\}$ represents the number of subtasks generated after the DNN task model is divided into blocks, its size depending on the initial task data size. Based on the size of the task data, we can categorize the task assignment modes into swarm assignment, partial assignment, and binary assignment. Swarm assignment indicates that the DNN task is processed serially through pipeline assignment across all wingman UAVs. Partial assignment means that the DNN task is processed serially through pipeline assignment among some wingman UAVs. Binary assignment entails that the DNN task is completely offloaded to a single wingman UAV for processing. The DNN task assignment mode ${M}$ is expressed as follows
@@
where $f(\cdot)$ represents the function that determines the task assignment mode (see Section VI). $\tau_{i}^{max}$ denotes the maximum latency allowed for the total completion time of each task, and $O$ represents the state space of the wingman UAV.
@@ III-C Communication Model
DNN task assignment is performed within the UAV swarm, and during the execution of the task, the UAV swarm maintains a fixed formation flying. At time slot $t$, the coordinates of UAV are designated as $\mathcal{P}^{n}=[x^{n}_{t},y^{n}_{t},z^{n}_{t}]$, and the distance between the two UAVs is defined as follows
@@
where $PL_{{FS},ref}(\mathfrak{f})$ represents the free space path loss per unit length, $\mathfrak{f}$ means the frequency of a radio wave is equal to the speed of light divided by the wavelength, $n_{CI}$ is the path loss index, and $\xi_{\sigma,CI}$ is the shadow fading parameter, which is typically assumed to follow a zero-mean Gaussian distribution.
@@
where $P_{n}$ is the information transmission power of the UAV, $P_{I}$ is the interference power, and $P_{N}$ is the noise power, denoted as $P_{N}=kBT$, where ${k}$ is the Boltzmann constant, ${B_{n}}$ is the bandwidth of the channel, and ${T}$ is the absolute temperature.
@@ III-D Mobile Model
Let $\boldsymbol{\varphi}_{q}$ denote the binary decision variable for the coordinate of the inspection target $\boldsymbol{g}_{\mathfrak{q}(q)}$; specifically, when the target point $\boldsymbol{g}_{\mathfrak{q}(q)}$ is inspected by the UAV, $\boldsymbol{\varphi}_{q}$ is equal to 1. Otherwise, $\boldsymbol{\varphi}_{q}$ is equal to 0. After the UAV swarm collects data from the target point $\boldsymbol{g}_{\mathfrak{q}(q)}$, it must process the data before reaching the next target point $\boldsymbol{g}_{\mathfrak{q}(q+1)}$. The Euclidean distance between the two target coordinates is denoted as $D^{\mathfrak{q}(q),\mathfrak{q}(q+1)}$. The total moving distance of the UAV to complete the inspection task is the sum of the distances of each segment along its flying path, as illustrated below
@@
Assuming that the UAV flies at a constant speed $\nu$, the time taken by the UAV swarm to travel from the current coordinate to the next coordinate is expressed as follows
@@
The total flying time for the process of the UAV executing the task is represented as follows
@@ III-E AoI Model
AoI defined in this paper aims to measure the duration from data collection to processing completion, incorporating waiting latency, transmission latency, and computation latency to reflect the real-time performance of DNN task execution. We assume that each UAV has a maximum memory capacity $m_{n}^{threshold}$, available energy $e_{n}^{threshold}$, and computing capacity $f_{n}$. The AoI is optimized by minimizing the UAV’s moving distance and refining the DNN task assignment strategy. The waiting latency for the data collected by the leader UAV to be processed by the DNN model is denoted as
@@
where $T_{i}$ represents the moment when the collected raw data is obtained on the leader UAV or when the subtask ${n}$ is generated on a wingman UAV. $T_{i^{\prime}}$ signifies the moment when the collected raw data is processed in the leader UAV or when the subtask ${i}$ is about to be sent from the wingman UAV ${n}$ to the wingman UAV ${n+1}$.
The latency for the intermediate result of the completed DNN task transmitted by wingman UAV ${n}$ is expressed as
@@
where $W_{i,l}$ is the data size output by the previous wingman UAV, specifically the data size before the split point $p_{i,n^{\prime}+1}$.
@@
Therefore, the execution latency of the entire DNN task ${AoI}$ and the overall latency $\mathcal{T}$ for the complete DNN task are as follows

### v3

@@ Abstract
Unmanned Aerial Vehicles (UAVs) offer high mobility and flexible deployment capabilities, making them ideal for Internet of Things (IoT) applications. However, the substantial amount of data generated by various applications within the existing low-altitude economic network requires processing through Deep Neural Networks (DNN) on UAVs, which is challenging due to their limited computational resources. To address this issue, we propose a two-stage optimization method for flight path planning and task allocation based on a mother-child UAV swarm system. In the first stage, we employ a greedy algorithm to solve the path planning problem by considering the task size of the target area to be inspected and the shortest flight path as constraints. The goal is to minimize both the flight path of the UAV and the overall cost of the system. In the second stage, we introduce a novel DNN task assignment algorithm that combines Multi-Agent Deep Deterministic Policy Gradient (MADDPG) and Generative Diffusion Models (GDM), named GDM-MADDPG. This algorithm takes advantage of the reverse denoising process of GDM to replace the actor network in MADDPG. It enables UAVs to generate specific DNN task assignment actions based on agents’ observations in a dynamic environment, thereby improving the efficiency of task scheduling and overall system performance. The simulation results demonstrate that our algorithm outperforms the benchmarks in terms of path planning, Age of Information (AoI), task completion, and system utility, demonstrating its effectiveness.
@@ I Introduction
Advancements in Unmanned Aerial Vehicles (UAVs) technology have led to their increasing use in various civil and military domains, including aerial object detection, rapid rescue operations in disaster-stricken areas [1, 2], emergency scenarios [3], and extensive agriculture [4]. Unlike traditional Internet of Things (IoT) devices [5], UAVs offer not only lower costs, but also superior mobility and enhanced data collection capabilities [6]. Although UAVs are well suited for the critical and complex tasks mentioned, managing the substantial data they generate remains a challenge [7]. With the rise of artificial intelligence (AI), deep neural networks (DNN) have been proposed to handle the extensive data produced [8]. Despite their efficiency in processing data, DNNs require significant resources due to their intricate structure and high performance demands. General methodologies segmented DNN models by layers, treating each layer as the primary computational unit, which often led to inefficiencies in collaborative inference due to the substantial computational demands of certain layers.
Given the size and cost constraints of UAVs, their power supply is often limited and computing capacity is relatively low. Generally, there are two primary approaches to tackle this issue. One approach involves deploying lightweight models on UAVs, which reduces computational demands and inference latency but sacrifices accuracy [9]. However, executing computation-intensive tasks independently with a single UAV presents significant challenges in achieving low latency and high inference accuracy. Another approach is to design from the perspective of the end-edge-cloud network architecture, enabling real-time, high-accuracy processing on UAVs with support from cloud and edge servers.
@@
Ref. Architecture Adaptive partitioning Path planning Latency Energy consumption Stability Age of information
@@ I-A Motivation and Challenges
Figure 2: The computational complexity, processing latency, and output data size of each layer of different models.
@@ I-B Summary of Contributions
  • We propose a mother-and-child UAV system that employs a high-altitude platform (HAP), such as an airship, equipped with multiple UAV launchers. From this HAP, several UAVs are launched and networked to support collaborative tasks. Then we design DNN tasks directly on UAVs, performing inference independently of the ground base station. This approach enhances the scalability of the system. Furthermore, we propose a novel greedy algorithm for UAV path planning that utilizes a fitness function to minimize flight distance, taking into account the varying sizes of tasks.
@@
II Related Works
In this section, we summarize the contributions of the related works and discuss the distinctions between prior research and our methodology.
@@ II-A Edge Intelligent for DNN Task
The partition of the DNN task represents a significant research challenge that involves dividing a DNN into several segments and distributing these segments between designated locations. The existing work provides valuable information to support the efficient DNN inference task. The authors in [24] present a joint design for task partitioning and offloading in DNN-enabled edge computing, involving a server and multiple devices. In [25], the authors propose a multi-task learning-based asynchronous advantage actor-critic approach, and explore distributed DNN inference through fine-grained model partitioning in edge computing. These studies overlook the mobility of UAV networks. The authors in [26] explore accelerating DNN-based task processing in vehicular edge computing through task partitioning and offloading. In [17], the authors propose an efficient allocation strategy using RL and investigate fast DNN inference in UAV swarms through multi-UAV collaboration. However, studies have not addressed the adaptation challenges associated with the partitioning of DNN tasks. In addition, the oversight of these studies may result in decreased network stability in certain resource-constrained networks due to suboptimal task assignment within the DNN.
@@ II-B MARL for Task Offloading
MARL enables each agent to optimize its policy through observations of both the environment and the policies of other agents [27, 28]. Many existing MARL-based approaches significantly improve task offload efficiency within dynamic networks [29, 30]. The authors in [31] propose a dedicated actor neural network for coordination and a scalable training algorithm for the offloading of UAV-aided MEC tasks. In [32], the authors present a DRL-based joint safe offloading and resource allocation scheme for vehicular edge computing networks. These studies have paid limited attention to this aspect. The authors in [33] present a mobility-aware service offloading and migration scheme with Lyapunov optimization and an MADDPG algorithm. The authors in [34] focus on optimizing task offloading and UAV trajectory under energy and queue latency constraints. However, most of these studies focus on optimizing network performance metrics, including task completion latency and energy consumption. Our research emphasizes DNN inference services, where the AoI for DNN tasks is a critical performance metric.
@@ II-C GAI for UAV Networks
GAI has the potential to effectively overcome the limitations of conventional AI and can be applied to optimize UAV networks, particularly to improve transmission rates, communication capacities, and energy efficiency [35, 36]. In [37], the authors present a radio propagation model with a conditional generative adversarial network. The authors in [38] introduce a two-stage generative neural network to predict link states and generate path losses, latencys, and arrival and departure angles in millimeter wave communication for UAVs. The authors in [39] explore the long-term optimization of joint task offloading and resource allocation in a multi-access edge computing network for UAVs. In [40], the authors present a hierarchical aerial computing system where UAVs collect tasks from ground devices and offload them to HAPs. Unlike the above schemes, our approach incorporates a path planning, GDM-based MARL decision-making framework specifically tailored for multi-UAV networks. This methodology aims to minimize the cost associated with DNN task completion while ensuring the system’s stability.
III System Model
In this section, we first present a novel network architecture, then introduce in detail the DNN task model, the communication model, the mobile model, the AoI model, and the energy consumption model, respectively.
@@ III-A Network Architecture
The network architecture is primarily composed of an HAP layer and a UAV swarm layer, as illustrated in Fig. 3. The HAP layer consists of a tethered platform with extensive computing and communication capabilities, such as an airship. This platform exchanges information with UAVs via wireless links to facilitate model training [41] and path planning (see Section V). The airship will transmit the final result of the path planning to the leader UAV. The UAV swarm comprises a leader UAV and multiple follow UAVs. All UAVs are isomorphic, with the total number of UAVs being ${n}=\left\{{1,2,\cdot\cdot\cdot,N}\right\}$, and each is equipped with trained DNN models. The leader UAV receives the results of the path planning transmitted by the airship and operates according to the predetermined route. It is also responsible for collecting and processing images and point cloud data of the target area, including tasks such as target detection and mapping. Thus, it functions as both a task producer and executor, as well as a decision maker for task assignment. The follow UAVs primarily execute the tasks assigned by the leader UAV, with the assignment of DNN tasks between follow UAVs being streamlined. The leader UAV evaluates the task size and the remaining resource status of the follow UAVs to determine whether to assign tasks. Therefore, to prevent task overload on the UAVs, extend the operational lifespan of the UAV swarm, and improve the timeliness of task processing and network stability, efficient task assignment decisions for the UAV swarm are particularly crucial.
@@ III-B DNN Task Model
The area where the task is performed in this paper consists of $q$ center coordinates of the target area, denoted as $q=1,2,...,\mathbb{W}$, and the sequence of the UAV inspection target coordinates is represented as $\mathfrak{q}(q)=[\mathfrak{q}(1),\ldots,\mathfrak{q}(\mathbb{W})]$. The data size in the target area is $\ell_{q}=\{1,...,\mathbb{L}_{\mathbb{W}}\}$. Assume that the number of DNN tasks in each synchronization cycle is $i=\{1,2,...,I\}$, the DNN type is $\kappa=\{1,...,\mathbb{K}\}$. The DNN model typically comprises $l$ layers, denoted as $l=\{1,2,...,L\}$. Each layer has a cache capacity $m_{i,l}$ and a computing requirement ${c}_{i,l}$, represented by a tuple $(m_{i,l},c_{i,l})$. $l_{\kappa}=\{1,2,...,L_{\kappa}\}$ represents a set of layers associated with a specific type of DNN task, including convolutional and pooling layers for feature extraction, as well as fully connected layers for classification [8]. The DNN task is decomposed into multiple subtasks for serial processing, and each follow UAV serves as an execution stage in this serial processing. This approach allows each subtask to offload and transfer intermediate results among different follow UAVs without requiring the processing of the entire task, thereby enhancing the computational efficiency and real-time performance of computationally intensive and latency-sensitive tasks [11].
In response to such practical needs, determining the optimal split point of the task is crucial in DNN task assignment. Assume that the split point of the ${i}$-th DNN task model is $P_{i,n^{\prime}}$, and $n^{\prime}\in\{1,2,...,N-1\}$ represents the number of subtasks generated after the DNN task model is divided into blocks, its size depending on the initial task data size. Based on the size of the task data, we can categorize the task assignment modes into swarm assignment, partial assignment, and binary assignment. Swarm assignment indicates that the DNN task is processed serially through pipeline assignment across all follow UAVs. Partial assignment means that the DNN task is processed serially through pipeline assignment among some follow UAVs. Binary assignment entails that the DNN task is completely offloaded to a single follow UAV for processing. The DNN task assignment mode ${M}$ is expressed as follows
@@
where $f(\cdot)$ represents the function that determines the task assignment mode (see Section VI). $\tau_{i}^{max}$ denotes the maximum latency allowed for the total completion time of each task, and $O$ represents the Observation space of the follow UAV.
@@ III-C Communication Model
DNN task assignment is performed within the UAV swarm, and during the execution of the task, the UAV swarm maintains a fixed formation flight. At time slot $t$, the coordinates of UAV are designated as $\mathcal{P}^{n}=[x^{n}_{t},y^{n}_{t},z^{n}_{t}]$, and the distance between the two UAVs is defined as follows
@@
where $PL_{{FS},ref}(\mathfrak{f})$ represents the free space path loss per unit length. $\mathfrak{f}$ means the frequency of a radio wave is equal to the speed of light divided by the wavelength. $n_{CI}$ is the path loss index, and $\xi_{\sigma,CI}$ is the shadow fading parameter, which is typically assumed to follow a zero-mean Gaussian distribution.
@@
where $P_{n}$ is the information transmission power of the UAV. $P_{I}$ is the interference power. $P_{N}$ is the noise power, denoted as $P_{N}=kBT$, where ${k}$ is the Boltzmann constant. ${B_{n}}$ is the channel bandwidth. ${T}$ is the absolute temperature.
@@ III-D Mobile Model
Let $\boldsymbol{\varphi}_{q}$ denote the binary decision variable for the coordinate of the inspection target $\boldsymbol{g}_{\mathfrak{q}(q)}$; specifically, when the target coordinate $\boldsymbol{g}_{\mathfrak{q}(q)}$ is inspected by the UAV, $\boldsymbol{\varphi}_{q}$ is equal to 1. Otherwise, $\boldsymbol{\varphi}_{q}$ is equal to 0. After the UAV swarm collects data from the target coordinate $\boldsymbol{g}_{\mathfrak{q}(q)}$, it must process the data before reaching the next target coordinate $\boldsymbol{g}_{\mathfrak{q}(q+1)}$. The Euclidean distance between the two target coordinates is denoted as $D^{\mathfrak{q}(q),\mathfrak{q}(q+1)}$. The total moving distance of the UAV to complete the inspection task is the sum of the distances of each segment along its flight path, as illustrated below
@@
Assuming that the UAV flies at a constant speed $\nu$, the time taken by the UAV swarm to travel from one coordinate to the next one is expressed as follows
@@
The total flight time for the process of the UAV executing the task is represented as follows
@@ III-E AoI Model
AoI defined in this paper aims to measure the duration from data collection to processing completion, incorporating waiting latency, transmission latency, and computation latency to reflect the real-time performance of DNN task assignment. We assume that each UAV has a maximum memory capacity $m_{n}^{threshold}$, available energy $e_{n}^{threshold}$, and computing capacity $f_{n}$. The AoI is optimized by minimizing the UAV’s moving distance and refining the DNN task assignment strategy. The waiting latency for the data collected by the leader UAV to be processed by the DNN model is denoted as
@@
where $T_{i}$ represents the moment when the collected raw data is obtained on the leader UAV or when the subtask ${n}$ is generated on a follow UAV. $T_{i^{\prime}}$ signifies the moment when the collected raw data is processed in the leader UAV or when the subtask ${i}$ is about to be sent from the follow UAV ${n}$ to the follow ${n+1}$.
The latency for the intermediate result of the completed DNN task transmitted by follow UAV ${n}$ is expressed as
@@
where $W_{i,l}$ is the data size output by the previous follow UAV, specifically the data size before the split point $p_{i,n^{\prime}+1}$.
@@
Therefore, the latency $\mathcal{T}$ for the complete DNN task and the execution latency of the entire DNN task ${AoI}$ are as follows